The 16 port patch panel has been installed in the rack and on the optics table. We have made 16, 24ft cables to go between these two patch panels. These cables are completed and installed. During testing we found one shorted cable and replaced it while strung up. 3 cables have a small piece of orange tape on them as a warning that the pin that makes the connected may be recessed too far in the cable. All 16 have been tested and are functioning however. Currently they are simply labelled according to numbers on the patch panel. We may update this labeling system. At minimum we will add these numbers to the cabel themselves in case they are unplugged for whatever reason.
Photo dump for the log. I have labelled all cables associated with the first patch panel. I have also cabeled (BNC only) the first filter cavity sled. See here.
Update to cabling for patch panels.